OpenStudy (anonymous):

I need help with this proportion problem please: Solve for x. 1/7=3/4x-1

OpenStudy (anonymous):

Flip the equation.

OpenStudy (anonymous):

How do I determine what x is?

OpenStudy (anonymous):

You need to flip the equation

OpenStudy (anonymous):

How do I flip the equation?

OpenStudy (anonymous):

Just put put the problem backwards like this: 3/4x−1=1/7

OpenStudy (anonymous):

Now we need to add 1 to both sides

OpenStudy (anonymous):

3/4x−1+1=1/7+1

OpenStudy (dbzfan836):

@SunStar28 There are easier ways to determine the answer. there was no need to flip.

OpenStudy (anonymous):

Well I am just following the way I solve it and how I like to solve it. :) I hope it isn't a problem

OpenStudy (anonymous):

3/4x=8/7

OpenStudy (dbzfan836):

Not a problem at all :D

OpenStudy (anonymous):

Multiply both sides by 4/3 :)

OpenStudy (anonymous):

Like this (4/3)*(3/4x)=(4/3)*(8/7)

OpenStudy (mathstudent55):

Cross multiply.

OpenStudy (dbzfan836):

That would cancel out our fraction next to the X

OpenStudy (mathstudent55):